Income Or Output Y Consumption Expenditure         C Investment Expenditure       I Government Expenditure          G...

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

$4,000

3,925

100

100

25

4,100

4,000

100

100

25

4,200

4,075

100

100

25

4,300

4,150

100

100

25

4,400

4,225

100

100

25

4,500

4,300

100

100

25

4,600

4,375

100

100

25

4,700

4,450

100

100

25

4,800

4,525

100

100

25

4,900

4,600

100

100

25

5,000

4,675

100

100

25

e) Suppose a tax cut of $50 is given with the information given in question1, what will be the new equilibrium level of income after the tax cut?
Answer #1

Answer : Before tax cut the equilibrium level of income is, Y = 4,400.

Because, Y = C + I + G - NX

=> Y = 4,225 + 100 + 100 - 25

=> Y = 4,400.

Now, if tax cut is $50 then this increase the government expenditure (G) by $50. So, the government expenditure will be, G = $150. As a result,

Y = 4,225 + 100 +150 -25

=> Y = 4,450

Therefore, after tax cut the new equilibrium income level is Y = 4,450.